OpenStudy (anonymous):

please help!! The weight W of an object varies inversely as the square of the distance d from the center of the earth. At sea level (3978 mi from the center of the earth), an astronaut weighs 124 lb. Find her weight when she is 123 mi above the surface of the earth and the spacecraft is not in motion. What is her weight ???

OpenStudy (anonymous):

inverse square relation : \[w=m \times \frac{ 1 }{ r^{2} }\] you can calculate the mass by comparing weight and r at sea level, then use the mass and 123 mi above to get the second weight.

OpenStudy (anonymous):

i still cant figure it out

OpenStudy (anonymous):

put into the equation the data that is given (first scenario, sea level): use the formula above

OpenStudy (anonymous):

first scenario: \[124 lb = m \frac{ 1 }{ r^{2} }\]

OpenStudy (anonymous):

do you know what is the distance from the center of the earth in the first scenario? r=

OpenStudy (anonymous):

3978

OpenStudy (anonymous):

and the weight was 124 lb, so we have: \[124lb=m \times \frac{1}{ 3978 ^{2} }\] mass is always the same, however, weight depends on gravity as well

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now we can solve for the mass by multiplying with 3978^2: \[124 \times 3978^{2} = m\]

OpenStudy (anonymous):

986544

OpenStudy (anonymous):

and then we look at the second scenario (above sea level): \[w = m \times \frac{ 1 }{ r^{2} }\] this time we are 123 mi ABOVE sea level, so 123 mi higher than before. the mass is the same (astronaut has not changed). formula: \[w = 986544 \times \frac{ 1 }{ (3978+123)^{2} }\]

OpenStudy (anonymous):

I got 116,6733

OpenStudy (anonymous):

lbs for the weight (with calculator)

OpenStudy (anonymous):

i didnt idk what i did wrong

OpenStudy (anonymous):

986544x4101^2

OpenStudy (anonymous):

i got 8091633888

OpenStudy (anonymous):

i need her in space

OpenStudy (anonymous):

the site was unresponsive yes you need her in space, that is why you add to the sea level another 123 mi

OpenStudy (anonymous):

we had \[124 \times 3978^{2}\] the result is 124 x 15824484 = 1962236016

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

and then you DIVIDE that by 1401^2

OpenStudy (anonymous):

ok then what do i do

OpenStudy (anonymous):

1962236016 / 4101^2

OpenStudy (anonymous):

700298.4

OpenStudy (anonymous):

the mass (1962236016) is the SAME, the distance from earth is now 4101 instead of 3978

OpenStudy (anonymous):

no, thats too much...

OpenStudy (anonymous):

mass (1962236016) divided by 1401^2=16818201

OpenStudy (anonymous):

weight = 1962236016/1401^2

OpenStudy (anonymous):

the numbers are very big because the distance from earth is so high

OpenStudy (anonymous):

from the center of the earth

OpenStudy (anonymous):

but 16818201 isnt her weight is it?

OpenStudy (anonymous):

no, 16818201 was just the new distance=1401 squared sorry if it's so confusing!

OpenStudy (anonymous):

lol its fine so whats next

OpenStudy (anonymous):

so that's what we divide the mass number 1962236016 with

OpenStudy (anonymous):

116.67

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

weight = 1962236016/1401^2 = 116.67 lbs when 123 mi above sea level

OpenStudy (anonymous):

rounded to the nearest tenth is 116.7??

OpenStudy (anonymous):

the weight decreased because earth has less power pulling her down when she already got away from the center more

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

116.7 isnt right

OpenStudy (anonymous):

when i typed it in it said it was wrong

OpenStudy (anonymous):

nvm thats right i was supossed to round to the hundreth thanks so mch

OpenStudy (anonymous):

then I am relieved ! :D your welcome